# Multivariate Variational Autoencoder

### v1

@@ V-C Ablation on Latent Dimensionality
Increasing $z$ improves reconstruction (MSE $\downarrow$) until saturation (e.g., CIFAR-10 stabilizes near 0.015–0.016 from $z{\geq}32$). Probe accuracy generally rises with $z$ before plateauing or slightly regressing at extreme sizes; clustering follows similar trends. Calibration is non-monotone: MVAE’s NLL/Brier/ECE typically improve from small to mid $z$ and fluctuate mildly at the largest $z$, reflecting bias–variance trade-offs in predictive entropy. Overall, MVAE’s advantage is most pronounced at mid-range $z$ (MNIST: 443232; CIFAR: 1616256256), where correlated posteriors deliver concurrent gains in calibration, linear separability, and clustering without sacrificing reconstruction; at very large $z$, both models converge in MSE and metric leadership may swap by dataset, while MVAE’s calibration remains robust.
@@
Figure 4: Latent-factor sweeps at $d_{z}=32$. MVAE exhibits smoother, monotone traversals (pose/expression) with fewer flat regions, indicating better factor geometry due to full-covariance posteriors.
@@ V-D Qualitative Evaluation
Figure 3 (first strip) shows triplets at $d_{z}{=}32$ (input, VAE recon, MVAE recon). Both models reconstruct faces sharply, yet MVAE preserves high-contrast facial contours (eye corners, nostrils, lip edges) with slightly reduced haloing around the cheeks and forehead. This is consistent with a posterior that allocates correlated uncertainty along semantically coherent directions, yielding crisper reconstructions without amplifying noise.
@@
We visualize $10\times 10$ traversals along two latent coordinates $(z_{i},z_{j})$ at $d_{z}{=}32$ for MVAE and VAE (Figure 4). Across the grid, both models produce smooth, monotone changes in pose/illumination. MVAE exhibits (i) slightly stronger local contrast (sharper eyelids, nasolabial fold, lip boundary), and (ii) clearer factor transitions (e.g., light direction versus subtle yaw/scale) with fewer degenerate tiles.
@@
The qualitative patterns align with the quantitative results reported in Section V-B: MVAE’s global coupling plus per-sample scales encourage correlated latent directions that track coherent face factors (lighting, micro-pose, soft tissue shading). This yields slightly sharper reconstructions and smoother, more axis-aligned traversals without sacrificing coverage of the latent plane.
@@ VI Discussion
MVAE targets a practical gap between diagonal-covariance VAEs and heavier posterior families by injecting structured correlations through a single, dataset-wide coupling matrix. Empirically, the approach offers three notable behaviors observed consistently across datasets and latent sizes: (i) calibration improves without sacrificing reconstruction, (ii) linear separability and clustering benefit most in the mid-$d_{z}$ regime, and (iii) latent-plane traversals become smoother and more axis-aligned, suggesting that correlated uncertainty aligns with semantically coherent factors (e.g., pose, illumination).
From a computational standpoint, the additional parameters in $\mathbf{C}\in\mathbb{R}^{d_{z}\times d_{z}}$ scale as $O(d_{z}^{2})$ but are amortized across samples; per-sample costs remain linear in $d_{z}$ due to the diagonal scales. In practice, this yields training/inference overheads that are modest relative to flow-augmented posteriors while retaining a closed-form KL. The qualitative and quantitative trends are strongest at mid-range latent sizes (MNIST: 443232; CIFAR: 1616256256), where diagonal posteriors underfit cross-dimensional structure and MVAE’s coupling brings the largest calibration and structure benefits. At very high $d_{z}$, both models converge in reconstruction and metric leadership may swap by dataset, but MVAE’s calibration typically remains competitive or superior.
@@ VII Conclusion
We introduced MVAE, a Gaussian, closed-form VAE with a full-covariance posterior realized via a global coupling matrix and per-sample scales. The method is simple to implement, analytically tractable, and empirically effective: it improves calibration and unsupervised structure with comparable or better reconstruction across diverse image regimes. Extensive experiments on Larochelle MNIST variants, Fashion-MNIST, CIFAR-10, and CIFAR-100, together with qualitative sweeps, support the claim that correlated posteriors yield more reliable and useful latent spaces than purely diagonal approximations. To promote reproducibility and future comparisons, we provide complete code, standardized splits/seeds, and metric pipelines.

### v2

@@ V-C Ablation on Latent Dimensionality
The effect of latent dimensionality ($z\in\{2,4,16,32,256,512\}$) exhibits a consistent qualitative trend across datasets: both models benefit from increased $z$ up to a point, but the relative advantages of MVAE tend to be most pronounced at low and moderate latent sizes.
At very low dimensions (z=2z{=}244), MVAE often compensates for the representational bottleneck by exploiting its structured covariance. On MNIST Basic and Fashion-MNIST, for example, MVAE achieves substantial relative improvements in accuracy and calibration at z=2z{=}2, and non-trivial gains in NMI/ARI (up to $\approx 5$1010% relative improvements) at z=16z{=}16 on MNIST Basic. On more challenging variants like MNIST Rotated or Background Random, MVAE still improves some metrics (e.g., Brier/ECE or NMI/ARI) at small $z$, but may sacrifice others (e.g., ARI on MNIST Rotated at z=2z{=}2), indicating that gains are dataset-dependent rather than uniform.
As $z$ increases, both VAE and MVAE approach similar reconstruction and classification performance, and differences become more subtle. MVAE tends to maintain slightly better or comparable calibration (Brier/ECE) and competitive clustering quality at intermediate sizes (z=16,32z{=}16,32), especially on MNIST Basic, MNIST Background Images, and CIFAR-10. At very high dimensions (z=256,512z{=}256,512), the diagonal VAE sometimes recovers small advantages in unsupervised metrics such as NMI and ARI on certain datasets, suggesting that the additional covariance flexibility is less critical once the latent space is highly overparameterized.
Overall, this dimensionality ablation suggests that MVAE’s multivariate posterior introduces a smoother and more robust trade-off between compactness and expressivity. Its benefits are strongest when capacity is limited or moderate, where correlated uncertainty provides a meaningful inductive bias, while remaining competitive with diagonal VAEs in overparameterized regimes.
@@
Figure 4: Latent-factor sweeps at $d_{z}=2$. MVAE exhibits smoother, monotone traversals (pose/expression) with fewer flat regions, indicating better factor geometry due to full-covariance posteriors.
@@ V-D Qualitative Evaluation
Figure 3 (first strip) illustrates qualitative reconstruction triplets at $d_{z}{=}2$, where each triplet corresponds to the original FreyFace input, the VAE reconstruction, and the MVAE reconstruction. Although both models capture the coarse facial structure, the MVAE produces noticeably sharper and more consistent reconstructions—particularly around the eyes and mouth regions—while avoiding the blurring typical of diagonal–covariance posteriors. This improvement stems from the MVAE’s ability to model correlated latent uncertainties through its coupling-matrix-based full covariance, enabling better preservation of subtle expression variations across frames. As latent dimensionality increases (not shown), both models improve reconstruction fidelity; however, the relative visual coherence and contrast balance often remain higher in MVAE samples, complementing the quantitative trends observed in Tables IVII.
@@
We visualize $10{\times}10$ latent traversals along two coordinates $(z_{i},z_{j})$ at $d_{z}{=}2$ for both models in Figure 4. Each grid samples a uniform mesh over the latent plane while keeping other dimensions fixed to zero. The VAE manifold (left) exhibits smooth but less semantically consistent transitions, often blending facial expressions and illumination changes. In contrast, the MVAE (right) produces a more structured and continuous latent surface, where horizontal and vertical directions correspond to interpretable variations such as pose and expression intensity. This qualitative smoothness supports the hypothesis that modeling correlated posteriors encourages globally coherent latent geometry—an effect that is broadly aligned with the improved clustering and calibration metrics reported in Section V-B.
@@
The FreyFace results collectively illustrate how introducing a multivariate latent structure reshapes both the statistical and geometric behavior of the learned representation. The standard VAE, constrained by a diagonal covariance, tends to encode redundant or axis-aligned directions, resulting in blurring and locally incoherent transitions in both reconstruction and latent traversal space. In contrast, the MVAE leverages sample-specific scales and a global coupling matrix to capture cross-dimensional correlations, effectively learning a smoother and more isotropic latent manifold. This manifests visually as sharper reconstructions and more interpretable traversals, and, on many datasets, as lower reconstruction error, improved calibration (lower Brier and ECE), and enhanced cluster consistency (higher NMI and ARI) at key latent sizes. Together, these findings suggest that structured uncertainty in the posterior acts as a useful inductive bias, helping align latent geometry with meaningful generative factors while maintaining compatibility with standard VAE training procedures.
@@ VI Discussion
The empirical findings across seven datasets highlight recurring advantages of the proposed Multivariate Variational Autoencoder (MVAE) over the standard VAE, while also revealing regimes where both models behave similarly. By extending the encoder to realize a full covariance via a global coupling matrix and per-sample scales, MVAE enriches the representational geometry of the latent space while retaining analytic tractability of the KL divergence. This modification yields both quantitative and qualitative effects that can be interpreted along three complementary dimensions.
First, representation quality. Across MNIST variants, Fashion-MNIST, and CIFAR-10/100, MVAE frequently produces sharper and more stable reconstructions, especially under complex background and texture conditions. The joint modeling of latent correlations enables the network to capture structured dependencies that diagonal posteriors cannot express, yielding smoother latent manifolds and more coherent generation at low and moderate latent dimensionalities. At very high dimensions, the two models often achieve comparable reconstruction performance, indicating that covariance structure is most beneficial when latent capacity is limited.
Second, calibration and uncertainty. In many settings, the inclusion of full covariance modeling results in improved calibration, reflected by lower Brier scores and ECE values for a wide range of $d_{z}$ on MNIST Basic, Fashion-MNIST, and CIFAR-10. On some datasets and latent sizes, MVAE matches but does not strictly dominate VAE, underscoring that the benefits of covariance modeling are dataset- and regime-dependent. Nevertheless, the overall pattern suggests that MVAE tends to learn uncertainty representations that more faithfully correspond to reconstruction confidence, narrowing the gap between generative modeling and reliability estimation.
Third, latent organization and clustering. Qualitative traversals (Figure 4) demonstrate that MVAE’s latent axes can encode interpretable directions of variation—pose, illumination, or expression—without explicit supervision. This structured latent geometry is echoed in clustering metrics (NMI and ARI), where MVAE often exhibits clear improvements at intermediate latent sizes, particularly on MNIST Basic and several corrupted MNIST variants. On some datasets and very high $z$, the diagonal VAE can recover slightly better clustering scores, indicating that covariance coupling is not a panacea but rather a strong inductive bias that interacts with dataset complexity and model capacity.
From a broader perspective, these results suggest that the independence assumption in conventional VAEs is not merely a convenient simplification but a limiting prior that can hinder posterior expressivity and calibration. By relaxing this assumption through a learnable, structured covariance with modest parameter overhead, MVAE offers a principled compromise between full-rank expressivity and computational efficiency, supporting both interpretability and performance across diverse visual domains.
@@ VII Conclusion
We have presented the Multivariate Variational Autoencoder (MVAE), a structured latent-variable model that generalizes the standard VAE by introducing correlated posteriors via a global coupling matrix and sample-specific scales. This design yields a full-covariance Gaussian posterior with a closed-form KL divergence, integrating seamlessly into standard VAE training pipelines.
Through extensive evaluation on seven datasets—ranging from binary MNIST variants to high-dimensional natural images—we find that MVAE often achieves superior or comparable performance across reconstruction, calibration, and clustering metrics, with the most pronounced gains appearing at low and moderate latent dimensionalities. While diagonal VAEs remain competitive, especially at very high latent dimensions and on certain unsupervised metrics, the empirical evidence indicates that even modest deviations from the diagonal assumption can provide meaningful representational benefits.
MVAE not only enhances reconstruction quality in challenging settings but also promotes smoother, semantically aligned latent manifolds and improved uncertainty estimates in many regimes. This supports the view that modeling multivariate uncertainty can act as an implicit regularizer, improving the geometric and probabilistic coherence of learned representations without substantial architectural complexity.
Future work will explore scaling the approach to hierarchical or multimodal settings, investigating interactions with expressive priors (e.g., VampPrior, flow-based posteriors), and embedding MVAE within self-supervised and contrastive learning frameworks. Overall, these findings establish MVAE as a tractable and interpretable extension of VAEs that unifies probabilistic rigor with enhanced latent-space structure, offering a practical path toward more geometrically grounded generative modeling.